

********************************************************************
UARS WEEKLY STATUS REPORT - 04 Jan 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues normal operations, taking solar data on all available
orbital tracking periods. We concur with SUSIM's request that the SSPP
be placed in closed loop solar pointing mode and have begun making only
solar observations.
***********
SUSIM
***********
SUSIM completed daily mid resolution and MgII solar scans during
periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continue. However, a hardware failure at GSFC caused an
interruption of DECNET communications from the Remote
Analysis Computers to the CMS for four days. To work
around this problem, several daily activity plans were
edited and entered manually on the CMS while others
were emailed to the MPG RAC via TCP/IP and then
transferred to the CMS when DECNET was restored. The
distribution of Sun Plans from the FOT to solar instruments
was also interrupted causing delays in the activity
plan preparation by those instruments.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 15 December 2001 (Orbit 56103, GMT Day 349)
through 21 December 2001 (Orbit 56206, GMT Day 355).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere
viewing) and is performing nominally under two-battery
operations. The instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 349/1236 56110 Power ON
PEM AXIS 351/0107 56133 Power ON (14 WATTS)
HALOE 354/2324 56192 Power OFF
The beta angle ranged between the angles of 30.94 degrees to
14.46 degrees for this report period. The beta angle is now
decreasing towards a minimum angle of 0 degrees on 25 December
2001 (DOY 359).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3
are currently maintaining greater than 25.5 V end-of-night (EON)
load bus voltage. Battery temperatures are stable with a
temperature delta between Battery 2 and Battery 3 of 0.66 to
1.31 degrees C. Additional battery performance data for each
battery on 08 December 2001 and 14 December 2001 are provided in
Appendix D.
The clock error ranged from
8.0 to 0.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the
initial excursion of ALERTS or Out-Of-Limits conditions detected
by the FOT for this report. A summary is listed below and the
details of ALERT and Out-Of-Limits occurrences are listed in
Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the
listed operational element(s) during this report period (see
Appendix A).
FOT = 0
HALOE = 2
PEM/AXIS = 1
UARS Anomaly Reports initiated during this report period are
listed in Appendix E and summarized below.
FOT = 2
One (1) Anomaly Report remained OPEN at the end of this report
period (see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 01/324 - 2287.498688
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period.
Anomaly Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF). This service should resume soon.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT conducted a test on 19 December 2001 with the new DSMC in
White Sands which is scheduled to replace NCC. The testing
included communication test messages, user performance data
and GCMRs for both SSA and MA supports. All configurations and
commands worked properly and no future tests are planned at this
time.
Friday, 21 December 2001 turned out to be a very successful day
for UARS. The Eclipse prototype system was successfully moved to
the Closed I/O net and established a link to the simulator
through the new TAC system. The first NOOP was sent to the
simulator and was verified by telemetry. Hundreds of NOOPs were
sent subsequent to the first NOOP, and all worked properly. This
was a very big step in the Eclipse transition process and the FOT
would like to thank all those involved in making this possible.
Also, on this day the FOT received word that the mini-CDHF was
online and should be ingesting UARS data again in order to
produce plots. After some minor print queue problems were
corrected the FOT was able to receive plots once again. This was
a very welcomed success after almost three months without any
trending capability.
MEETINGS
The FOT attended a UARS Eclipse Weekly on Monday, December 17,
2001.
The FOT attended a UARS Eclipse Working Meeting on Thursday,
December 20, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 24 December 2001 (DOY
358). This maneuver will be a Forward to Reverse Flight maneuver
(Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 21 December 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
56110 349/1236 01-728 HALOE Power ON
56133 351/0107 01-729 PEM AXIS Power ON (14 WATTS)
56192 354/2324 01-730 HALOE Power OFF
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 21 December 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 21 December 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
-None this report period.
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 21 December 2001
15 December 2001 - GMT Day 349 Beta = 30.94 deg, SA Pos =
269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ------- ------- --- ----- ------
1 -560.0/-728.0 -3.16/-4.38 +0.0/-0.0 10.72 0.000 0.00
2 +000.0/-504.0 +5.47/+4.48 +28.4/-12.8 25.72 1.040 18.76
3 +224.0/-112.0 +4.48/+3.82 +27.2/-12.8 25.56 1.040 18.41
21 December 2001 - GMT Day 355 Beta = 14.46 deg, SA Pos =
269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ------- ------- ---- ----- ------
1 -560.0/-728.0 -3.47/-4.38 +0.0/-0.0 10.72 0.000 0.00
2 +000.0/-224.0 +4.48/+3.82 +28.8/-12.8 26.40 1.020 19.34
3 +224.0/-224.0 +3.17/+2.85 +28.0/-12.8 26.40 1.020 18.87
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 21 December 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-083 FOT 56157 352/1551 24145 Patch Panel/TAC 5
01-084 FOT 56174 353/1913 N/A New TAC 4 Reboot
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-083 FOT 56157 352/1551 24145 Patch Panel/TAC 5
01-084 FOT 56174 353/1913 N/A New TAC 4 Reboot
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-75 FOT see list Dec 01 14170 Generic Late Acquisition
***********************************************************
REPORTING PERIOD
This report covers 22 December 2001 (Orbit 56207, GMT Day 356)
through 28 December 2001 (Orbit 56311, GMT Day 362).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere
viewing) and is performing nominally under two-battery
operations. The instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
---------- ---- ----- -------
PEM 356/2248 56221 POWER OFF
SUSIM 357/2137 56236 POWER OFF
PEM AXIS 357/2137 56236 POWER OFF
ACRIM 360/1641 56277 POWER OFF/ON
SUSIM 361/1318 56290 POWER ON (Begin)
PEM 362/0038 56297 ZEPS/HEPS Power On - Day Only
SUSIM 362/1317 56305 Power On - Complete
A Forward-to-Reverse Yaw Around maneuver (number 105) was
successfully executed in Orbit 56250 beginning 358/18:45:19 and
completing at 358/19:22:29 on 24 December 2001. The SOLSTICE
instrument was in day only and all other instruments were in
their off or standby mode (including HRDI and PEM). Due to the
SA position at 269 degrees, the effective night length of the
spacecraft after the maneuver was approximately 57.1 minutes.
Post maneuver analysis of data indicate that during the Yaw
Around, the end-of-night load bus voltage was 24.9 volts and the
average depths of discharge (DODs) on Batteries 2 and 3 were
approximately 20.16 and 19.81 percent respectively. The next
(forward-to-reverse) YAW is scheduled for 03 February 2002 (DOY
034).
The beta angle ranged between the angles of 30.94 degrees to
09.68 degrees for this report period. The beta angle is now
decreasing towards a minimum angle of 0 degrees on 03 February
2002 (DOY 034).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3
are currently maintaining greater than 26.4 V end-of-night (EON)
load bus voltage. Battery temperatures are stable with a
temperature delta between Battery 2 and Battery 3 of 0.97 to 1.31
degrees C. Additional battery performance data for each battery
on 22 December 2001 and 28 December 2001 are provided in Appendix
D.
The clock error ranged from -8.0 to 0.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the
initial excursion of ALERTS or Out-Of-Limits conditions detected
by the FOT for this report. A summary is listed below and the
details of ALERT and Out-Of-Limits occurrences are listed in
Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the
listed operational element(s) during this report period (see
Appendix A).
FOT = 6
PEM/AXIS = 3
SOLSTICE = 2
HRDI = 1
SUSIM = 3
ACRIM = 1
UARS Anomaly Reports initiated during this report period are
listed in Appendix E and summarized below.
FOT = 0
One (1) Anomaly Report remained OPEN at the end of this report
period (see Appendix E).
TRANSPONDER FREQUENCIES:
* Transponder A: 01/333 - 2287.496855
* Transponder B: 01/324 - 2287.498688
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period.
Anomaly Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF). This service should resume soon.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Another successful yaw maneuver was executed on 24 December 2001.
The TAM and FSS configuration continues to work well at
maintaining attitude. The first FSS residuals after the yaw
maneuver were approximately 400 arc seconds.
The FOT continued its testing with the new TAC systems this week.
The FOT in conjunction with the TAC developers are still trying
hard to resolve the open issues on the TAC system. Several more
tests were run in an attempt to isolate the problem without
success. Only a few tests remain that could give any further
insight into the problem. These tests will be run next week.
The UARS FOT would also like to wish everyone a Happy New Year
and say thank you for the much appreciated hard work and
determination that everyone has given.
MEETINGS
The FOT attended a UARS Eclipse Working Meeting on Thursday,
December 27, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 03 February 2002 (DOY
034). This maneuver will be a Reverse to Forward Flight maneuver
(Southern Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 28 December 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
56219 356/1950 01-731 OBC/CLK Clk Adj 0.0 to -20.6 msec/day
56221 356/2248 01-732 PEM POWER OFF
56236 357/2137 01-733 SOLSTICE TO DAY ONLY OPS
56236 357/2137 01-734 HRDI TO LOW PWR
56236 357/2137 01-735 SUSIM POWER OFF
56236 357/2137 01-736 PEM AXIS POWER OFF
56237 358/0026 01-737 OBC/CLK Clk Adj -20.6 to 0.0 msec/day
56248 358/1807 01-738 ACS YAW SCRIPT - BEGIN
56248 358/1807 01-739 ACS/TAM TAM UPDATE MODE DISABLE
56250 358/2005 01-738 ACS YAW SCRIPT - END
56250 358/2005 01-740 ACS/TAM TAM UPDATE MODE ENABLE
56277 360/1641 01-741 ACRIM POWER OFF/ON
56282 361/0039 01-742 SOLSTICE TO FULL TIME OPS
56290 361/1318 01-744 SUSIM POWER ON (Begin)
56297 362/0038 01-743 PEM ZEPS/HEPS Power On - Day Only
56305 362/1317 01-744 SUSIM Power On - Complete
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 28 December 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 28 December 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
56219 356/1950 Clk Adj 0.0 to -20.6 msec/day
56237 358/0026 Clk Adj -20.6 to 0.0 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 28 December 2001
22 December 2001 - GMT Day 356 Beta = 30.94 deg, SA Pos =
269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ----- -----
1 -560.0/-728.0 -3.47/-4.38 +0.0/ -0.0 10.72 0.000 0.00
2 +000.0/-224.0 +4.48/+3.82 +28.8/-12.8 26.40 1.020 19.34
3 +224.0/-224.0 +3.17/+2.85 +28.0/-12.8 26.40 1.020 18.87
28 December 2001 - GMT Day 362 Beta = 09.68 deg, SA Pos =
269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ----- -----
1 -560.0/-728.0 -4.08/-4.99 +0.0/ -0.0 10.72 0.000 0.00
2 +056.0/-448.0 +4.48/+3.82 +28.0/-12.8 26.08 1.010 20.16
3 +224.0/-112.0 +3.17/+2.52 +27.2/-12.8 26.08 1.010 19.81
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 28 December 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-75 FOT see list Dec 01 14170 Generic Late Acquisition